### v2

@@ 1 Introduction
To some extent, the increasing performance of deep restoration models largely stems from the increasing network receptive field. First, a large receptive field allows the network to capture information from a wider region, enabling it to refer to more pixels to facilitate the reconstruction of the anchor pixel. Second, with a larger receptive field, the restoration network can extract higher-level patterns and structures in the image, which can be crucial for some structure preservation tasks such as image denoising. Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent art [9] also points out that activating more pixels usually leads to better restoration results.
@@
Figure 1: The Effective Receptive Field (ERF) visualization [40] for EDSR [37], RCAN [76], SwinIR [36], HAT [9], and the proposed MambaIR. A larger ERF is indicated by a more extensively distributed dark area. Only the proposed MambaIR achieves a significant global effective receptive field.
@@
Despite possessing many attractive properties, it appears that there exists an inherent choice dilemma between global receptive fields and efficient computation for current image restoration backbones. For CNN-based restoration networks [37, 76], although the effective receptive field is limited (as shown in Fig. 1(a)), it is appropriate for resource-constrained device deployments due to the favorable efficiency of convolution parallel operations. By contrast, Transformer-based image restoration methods usually set the number of tokens to the image resolution [36, 8, 9], therefore, despite the global receptive field, directly using the standard Transformer [57] will come at an unacceptable quadratic computational complexity. Moreover, employing some efficient attention techniques such as shifted window attention [39] for image restoration, usually comes at the expense of a globally effective receptive field (as shown in  Fig. 1(b)), and does not intrinsically escape out of the trade-off between a global receptive field and efficient computation.
@@
Recently, structured state-space sequence models (S4), especially the improved version Mamba, have emerged as an efficient and effective backbone for constructing deep networks [21, 55, 17, 46, 19]. This development hints at a potential resolution to balancing global receptive field and computational efficiency in image restoration. In detail, the discretized state space equations in Mamba can be formalized into a recursive form and can model very long-range dependencies when equipped with specially designed structured reparameterization [20]. This means that Mamba-based restoration networks can naturally activate more pixels, thus improving the reconstruction quality. Furthermore, the parallel scan algorithm [19] renders Mamba to process each token in a parallel fashion, facilitating efficient training on modern hardware such as GPU. The above promising properties motivate us to explore the potential of Mamba to achieve efficient long-range modeling for image restoration networks.
@@
  • We are the first to work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
@@ 2.1 Image Restoration
Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [14] for image super-resolution, DnCNN [70] for image denoising, ARCNN [13] for JPEG compression artifact reduction, etc. Early attempts usually elaborate CNNs with techniques such as residual connection [29, 67, 6], dense connection [60, 78] and others [31, 62, 18, 76, 12] to improve model representation ability. Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies. As Transformers have been proven to be a strong competitor for CNNs in multiple computer vision tasks [5, 15, 39], using Transformers for image restoration appears promising. Despite the global receptive field, it still faces specific challenges from the quadratic computational complexity of the self-attention mechanism [57]. To address this, IPT [8] divides one image into several small patches and processes each patch independently with self-attention. SwinIR [36] further introduces shifted window attention [39] to improve the performance. In addition, progress continues to be made in designing efficient attention mechanisms for restoration [66, 9, 32, 11, 63, 10]. Nonetheless, efficient attention design usually comes at the expense of global effective receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
@@ 2.2 State Space Models
State Space Models (SSMs) [22, 21, 55], stemming from classics control theory [28], are recently introduced to deep learning as a competitive backbone for state space transforming. The promising property of linearly scaling with sequence length in long-range dependency modeling has attracted great interest from searchers. For example, the Structured State-Space Sequence model (S4) [21] is a pioneer work for the deep state-space model in modeling the long-range dependency. Later, S5 layer [55] is proposed based on S4 and introduces MIMO SSM and efficient parallel scan. Moreover, H3 [17] achieves promising results that nearly fill the performance gap between SSMs and Transformers in natural language.  [46] further improve S4 with gating units to obtain the Gated State Space layer to boost the capability. More recently, Mamba [19], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length. Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [38, 81], video understanding [58], biomedical image segmentation [42] and others [26, 50]. In this work, we explore the potential of Mamba to image restoration with restoration-specific designs to serve as a simple but effective baseline for future work.
@@ 4.1 Experimental Settings
Dataset and Evaluation. Following the setup in previous works [36, 66], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, and real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising). We employ DIV2K [56] and Flickr2K [37] to train classic SR models and use only DIV2K to train lightweight SR models. Moreover, we use Set5 [4], Set14 [65], B100 [44], Urban100 [24], and Manga109 [45] to evaluate the effectiveness of different SR methods. For gaussian color image denoising, we utilize DIV2K [56], Flickr2K [37], BSD500 [3], and WED [43] as our training datasets. Our testing datasets for guassian color image denoising includes BSD68 [44], Kodak24 [16], McMaster [74], and Urban100 [24]. For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [52] dataset for testing. Following [36, 76], we denote the model as MambaIR+ when self-ensemble strategy [37] is used in testing. The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space. In addition, we also include other restoration tasks such as JPEG compression artifact reduction, more experimental results can be seen in the supplementary material.
Table 1: Ablation experiments for different design choices of RSSB.
settings #param MACs Set5 Set14 B100 Urban100 Manga109
(0)baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
(1)+w/o Conv 11.9M 316G 38.48 34.54 32.56 34.04 40.20
(2)+w/o Conv+CA 11.8M 315G 38.55 34.64 32.57 34.06 40.14
(3)+replace with MLP 14.3M 379G 38.55 34.68 32.59 34.22 40.13
@@ 4.2 Ablation Study
Effects of different designs of RSSB. As the core component, the RSSB can improve Mamba with restoration-specific priors. In this section, we ablate different components of the RSSB. The results, presented in Tab. 1, indicate that (1) directly using the standard Mamba to process flattened sequences can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction. (2) Without using additional convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results, which also supports our previous analysis. (3) Replacing Conv+ChanelAttention with MLP, whose resulted structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, the behavior of these two modules is different and thus accustomed block structure should be considered for further improvements.
@@
Table 2: Ablation experiments for different scan modes in VSSM.
scan mode #param MACs Set5 Set14 B100 Urban100 Manga109
one-direction 16.7M 439G 38.53 34.63 32.58 34.06 40.31
two-direction 16.7M 439G 38.56 34.60 32.56 33.96 40.14
baseline 16.7M 439G 38.57 34.67 32.58 34.15 40.28
@@
MambaIR+ ×$\times 4$ 31.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
Effects of Different Scan Modes in VSSM. To allow Mamba to process 2D images, the feature map needs to be flattened before being iterated by the state-space equation. Therefore, the unfolding strategy is particularly important. In this work, we follow  [38] which uses scans in four different directions to generate scanned sequences. Here, we ablate different scan modes to study the effects, the results are shown in Tab. 2. Compared with one-direction(top-left to bottom-right) and two-direction(top-left to bottom-right, bottom-right to top-left), using four directions of scanning allows the anchor pixel to perceive a wider range of neighborhoods, thus achieving better results. Surprisingly, using more scanning directions brings an almost negligible increase in the number of parameters and computational complexity. This benefit may facilitate future exploration of more effective scanning mechanisms for the Mamba-based image restoration networks. We also include other ablation experiments, such as the layer number of RSSBs, please see supplementary material for more analysis.
@@ 4.3 Comparison on Image Super-Resolution
Classic Image Super-Resolution. Tab. 3 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods. Thanks to the significant global receptive field, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors. For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration. We also give visual comparisons in Fig. 4, and it can be seen that our method can facilitate the reconstruction of sharp edges and natural textures.
Table 4: Model size comparisons. We compare the complexity on the ×$\times 4$ classic SR model of different methods, the output size is set to $3\times 640\times 640$ for MACs.
Method EDSR [37] RCAN  [76] SRFBN  [35] HAN  [51] CSNLN  [48] SwinIR  [36] MambaIR (Ours)
Params (M) 43.1 15.6 3.6 16.1 7.2 11.9 16.7
Mult-Adds $(\mathrm{G})$ 1,286 407 498 420 103,640 336 439
PSNR on Urban100 (dB) 26.64 26.82 26.60 26.85 27.22 27.45 27.68
PSNR on Manga109 (dB) 31.02 31.22 31.15 31.42 31.43 32.03 32.32
@@
Model Complexity Comparison.  Tab. 4 provides the comparison of different methods on parameters and MACs. Our MambaIR has comparable parameters and MACs compared with SwinIR, but can achieve a global receptive field and better PSNR performance. We further give comparison results of computational complexity with input sizes in  Fig. 5. As one can see, our method is far more efficient than the full-attention baseline [57] and exhibits linear complexity with input resolution which is similar to the efficient attention techniques such as SwinIR. These observations above suggest that out MambaIR has similar scale properties as shifted window attention, while possessing a global receptive field similar to standard full attention.
Lightweight Image Super-Resolution. To demonstrate the scalability of our method, we train the Mamba-light model and compare it with state-of-the-art lightweight image SR methods. Following previous works [80, 41], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution). Tab. 5 shows the results. It can be seen that our MambaIR-light outperforms SwinIR-light [36] by up to 0.17dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs. The performance results demonstrate the scalability and efficiency of our method.
@@
LatticeNet [41] $\times$2 756K 169.5G 38.15 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
SwinIR-light [36] $\times$2 878K 195.6G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
SRFormer-light [80] $\times$2 853K 236G 38.23 0.9613 33.94 0.9209 32.36 0.9019 32.91 0.9353 39.28 0.9785
Ours $\times$2 859K 198.1G 38.16 0.9610 34.00 0.9212 32.34 0.9017 32.92 0.9356 39.31 0.9779
@@
SwinIR-light [36] $\times$3 886K 87.2G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
SRFormer-light [80] $\times$3 861K 105G 34.67 0.9296 30.57 0.8469 29.26 0.8099 28.81 0.8655 34.19 0.9489
Ours $\times$3 867K 88.7G 34.72 0.9296 30.63 0.8475 29.29 0.8099 29.00 0.8689 34.39 0.9495
@@
SwinIR-light [36] $\times$4 897K 49.6G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
SRFormer-light [80] $\times$4 873K 62.8G 32.51 0.8988 28.82 0.7872 27.73 0.7422 26.67 0.8032 31.17 0.9165
Ours $\times$4 879K 50.6G 32.51 0.8993 28.85 0.7876 27.75 0.7423 26.75 0.8051 31.26 0.9175
@@ 4.4 Comparison on Image Denoising
Gaussian Color Image Denoising. The results of gaussian color image denoising are shown in  Tab. 6. Following [67, 70], the compared noise levels include 15, 25 and 50. As one can see, our model achieves better performance than all compared methods. In particular, it surpasses the SwinIR [36] by up to 0.25dB with $\sigma$=50 on the Urban100 dataset. We also give a visual comparison in Fig. 7. Thanks to the global receptive field, our MambaIR can achieve better structure preservation, leading to clearer edges and natural shapes.
@@
Real Image Denoising. We further turn to the real image denoising task to evaluate the robustness of MambaIR when facing real-world degradation. Following [63], we adopt the progressive training strategy for fair comparison. The results, shown in Tab. 7, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [61] and outperforms other methods such as Uformer [61] by 0.12dB PSNR on SIDD dataset. The promising results indicate the ability of our method in real image denoising.
Table 7: Quantitative comparison on real image denosing with state-of-the-art methods.

### v3

@@ 1 Introduction
To some extent, the increasing performance of deep restoration models largely stems from the increasing network receptive field. First, a large receptive field allows the network to capture information from a wider region, enabling it to refer to more pixels to facilitate the reconstruction of the anchor pixel. Second, with a larger receptive field, the restoration network can extract higher-level patterns and structures in the image, which can be crucial for some structure preservation tasks such as image denoising. Finally, Transformer-based restoration methods which possess larger receptive fields experimentally outperform CNN-based methods, and the recent work [10] also points out that activating more pixels usually leads to better restoration results.
@@
Figure 1: The Effective Receptive Field (ERF) visualization [46, 14] for EDSR [42], RCAN [88], SwinIR [41], HAT [10], and the proposed MambaIR. A larger ERF is indicated by a more extensively distributed dark area. The proposed MambaIR achieves a significant global effective receptive field.
@@
Despite possessing many attractive properties, it appears that there exists an inherent choice dilemma between global receptive fields and efficient computation for current image restoration backbones. For CNN-based restoration networks [42, 89], although the effective receptive field is limited (as shown in Fig. 1(a)), it is appropriate for resource-constrained device deployments due to the favorable efficiency of convolution parallel operations. By contrast, the Transformer-based image restoration methods usually set the number of tokens to the image resolution [41, 8, 10], therefore, despite the global receptive field, directly using the standard Transformer [65] will come at an unacceptable quadratic computational complexity. Moreover, employing some efficient attention techniques such as shifted window attention [45] for image restoration, usually comes at the expense of a globally effective receptive field (as shown in  Fig. 1(b)), and does not intrinsically escape out of the trade-off between a global receptive field and efficient computation.
@@
Recently, structured state-space sequence models (S4), especially the improved version Mamba, have emerged as an efficient and effective backbone for constructing deep networks [24, 62, 18, 52, 22]. This development hints at a potential solution to balancing global receptive field and computational efficiency in image restoration. In detail, the discretized state space equations in Mamba can be formalized into a recursive form and can model very long-range dependencies when equipped with specially designed structured reparameterization [23]. This means that Mamba-based restoration networks can naturally activate more pixels, thus improving the reconstruction quality. Furthermore, the parallel scan algorithm [22] renders Mamba to process each token in a parallel fashion, facilitating efficient training on modern hardware such as GPU. The above promising properties motivate us to explore the potential of Mamba to achieve efficient long-range modeling for image restoration networks.
@@
  • We are the first work to adapt state space models for low-level image restoration via extensive experiments to formulate MambaIR, which acts as a simple but effective alternative for CNN- and Transformer-based methods.
@@ 2.1 Image Restoration
Image restoration has been significantly advanced since the introduction of deep learning by several pioneering works, such as SRCNN [16] for image super-resolution, DnCNN [81] for image denoising, ARCNN [15] for JPEG compression artifact reduction, etc. Early attempts usually elaborate CNNs with techniques such as residual connection [34, 6], dense connection [68, 89] and others [36, 70, 19, 13] to improve model representation ability. Despite the success, CNN-based restoration methods typically face challenges in effectively modeling global dependencies. As transformer have proven its effectiveness in multiple tasks, such as time series [43], 3D cloud [77, 75], and multi-modal [20, 4, 21, 86], using transformer for image restoration appears promising. Despite the global receptive field, transformer still faces specific challenges from the quadratic computational complexity of the self-attention [65]. To address this, IPT [8] divides one image into several small patches and processes each patch independently with self-attention. SwinIR [41] further introduces shifted window attention [45] to improve the performance. In addition, progress continues to be made in designing efficient attention for restoration [78, 10, 38, 12, 72, 11, 63, 26, 85, 9]. Nonetheless, efficient attention design usually comes at the expense of global receptive fields, and the dilemma of the trade-off between efficient computation and global modeling is not essentially resolved.
@@ 2.2 State Space Models
State Space Models (SSMs) [25, 24, 62], stemming from classics control theory [33], are recently introduced to deep learning as a competitive backbone for state space transforming. The promising property of linearly scaling with sequence length in long-range dependency modeling has attracted great interest from searchers. For example, the Structured State-Space Sequence model (S4) [24] is a pioneer work for the deep state-space model in modeling the long-range dependency. Later, S5 layer [62] is proposed based on S4 and introduces MIMO SSM and efficient parallel scan. Moreover, H3 [18] achieves promising results that nearly fill the performance gap between SSMs and Transformers in natural language.  [52] further improve S4 with gating units to obtain the Gated State Space layer to boost the capability. More recently, Mamba [22], a data-dependent SSM with selective mechanism and efficient hardware design, outperforms Transformers on natural language and enjoys linear scaling with input length. Moreover, there are also pioneering works that adopt Mamba to vision tasks such as image classification [44, 92], video understanding [66, 37], biomedical image segmentation [48, 71] and others [31, 56, 28, 76, 59]. In this work, we explore the potential of Mamba to image restoration with restoration-specific designs to serve as a simple but effective baseline for future work.
@@ 4.1 Experimental Settings
Dataset and Evaluation. Following the setup in previous works [41, 78], we conduct experiments on various image restoration tasks, including image super-resolution (i.e., classic SR, lightweight SR, real SR) and image denoising (i.e., Gaussian color image denoising and real-world denoising), and JPEG compression artifact reduction (JPEG CAR). We employ DIV2K [64] and Flickr2K [42] to train classic SR models and use DIV2K only to train lightweight SR models. Moreover, we use Set5 [5], Set14 [74], B100 [50], Urban100 [29], and Manga109 [51] to evaluate the effectiveness of different SR methods. For gaussian color image denoising, we utilize DIV2K [64], Flickr2K [42], BSD500 [3], and WED [49] as our training datasets. Our testing datasets for guassian color image denoising includes BSD68 [50], Kodak24 [17], McMaster [84], and Urban100 [29]. For real image denoising, we train our model with 320 high-resolution images from SIDD [1] datasets, and use the SIDD test set and DND  [58] dataset for testing. Following [41, 89], we denote the model as MambaIR+ when self-ensemble strategy [42] is used in testing. The performance is evaluated using PSNR and SSIM on the Y channel from the YCbCr color space. Due to page limit, the results of JPEG CAR are shown in the supplementary material.
@@ 4.2 Ablation Study
Effects of different designs of RSSB. As the core component, the RSSB can improve Mamba with restoration-specific priors. In this section, we ablate different components of the RSSB. The results, presented in Tab. 1, indicate that (1) applying 1D scan on flattened images can lead to local pixel forgetting, and the utilization of simple convolution layers can effectively enhance the local interaction. (2) Without using additional convolution and channel attention, i.e., directly employing off-the-shelf Mamba for restoration, can only obtain sub-optimal results, which also supports our previous analysis. (3) Replacing Conv+ChanelAttention with MLP, whose resulted structure will be similar to Transformer, also leads to unfavorable results, indicating that although both SSMs and Attention have the global modeling ability, the behavior of these two modules is different and thus accustomed block structure should be considered for further improvements.
Table 1: Ablation experiments for different design choices of RSSB.
settings Set5 Set14 Urban100
(1)remove Conv 38.48 34.54 34.04
(2)remove Conv+CA 38.55 34.64 34.06
(3)replace with MLP 38.55 34.68 34.22
Table 2: Ablation experiments for different scan modes in VSSM.
scan mode Set5 Set14 Urban100
one-direction 38.53 34.63 34.06
two-direction 38.56 34.60 33.96
baseline 38.57 34.67 34.15
@@
MambaIR+ ×$\times 4$ 33.13 0.9054 29.25 0.7971 28.01 0.7510 27.80 0.8303 32.48 0.9281
Effects of Different Scan Modes in VSSM. To allow Mamba to process 2D images, the feature map needs to be flattened before being iterated by the state-space equation. Therefore, the unfolding strategy is particularly important. In this work, we follow [44] which uses scans in four different directions to generate scanned sequences. Here, we ablate different scan modes to study the effects, the results are shown in Tab. 2. Compared with one-direction (top-left to bottom-right) and two-direction (top-left to bottom-right, bottom-right to top-left), using four directions of scanning allows the anchor pixel to perceive a wider range of neighborhoods, thus achieving better results. We also include other ablation experiments, such as the layer number of RSSBs, please see supplementary material for more analysis.
@@ 4.3 Comparison on Image Super-Resolution
Classic Image Super-Resolution. Tab. 3 shows the quantitative results between MambaIR and state-of-the-art super-resolution methods. Thanks to the significant global receptive field, our proposed MambaIR achieves the best performance on almost all five benchmark datasets for all scale factors. For example, our Mamba-based baseline outperforms the Transformer-based benchmark model SwinIR by 0.41dB on Manga109 for ×$\times 2$ scale, demonstrating the prospect of Mamba for image restoration. We also give visual comparisons in Fig. 4, and our method can facilitate the reconstruction of sharp edges and natural textures.
@@
Model Complexity Comparison. We give comparison results of computational complexity with varying input sizes in  Fig. 5. As one can see, our method is far more efficient than the full-attention baseline [65] and exhibits linear complexity with input resolution which is similar to the efficient attention techniques such as SwinIR. These observations above suggest that out MambaIR has similar scale properties as shifted window attention, while possessing a global receptive field similar to standard full attention.
Lightweight Image Super-Resolution. To demonstrate the scalability of our method, we train the MambaIR-light model and compare it with state-of-the-art lightweight image SR methods. Following previous works [91, 47], we also report the number of parameters (#param) and MACs (upscaling a low-resolution image to $1280\times 720$ resolution). Tab. 4 shows the results. It can be seen that our MambaIR-light outperforms SwinIR-light [41] by up to 0.34dB PSNR on the ×$\times 4$ scale Manga109 dataset with similar parameters and MACs. The performance results demonstrate the scalability and efficiency of our method.
@@
LatticeNet [47] $\times$2 756K 169.5G 38.13 0.9610 33.78 0.9193 32.25 0.9005 32.43 0.9302 - -
SwinIR-light [41] $\times$2 910K 122.2G 38.14 0.9611 33.86 0.9206 32.31 0.9012 32.76 0.9340 39.12 0.9783
Ours $\times$2 905K 167.1G 38.13 0.9610 33.95 0.9208 32.31 0.9013 32.85 0.9349 39.20 0.9782
@@
SwinIR-light [41] $\times$3 918K 55.4G 34.62 0.9289 30.54 0.8463 29.20 0.8082 28.66 0.8624 33.98 0.9478
Ours $\times$3 913K 74.5G 34.63 0.9288 30.54 0.8459 29.23 0.8084 28.70 0.8631 34.12 0.9479
@@
SwinIR-light [41] $\times$4 930K 31.8G 32.44 0.8976 28.77 0.7858 27.69 0.7406 26.47 0.7980 30.92 0.9151
Ours $\times$4 924K 42.3G 32.42 0.8977 28.74 0.7847 27.68 0.7400 26.52 0.7983 30.94 0.9135
@@ 4.4 Comparison on Image Denoising
Gaussian Color Image Denoising. The results of gaussian color image denoising are shown in  Tab. 5. Following [81, 79], the compared noise levels include 15, 25 and 50. As one can see, our model achieves the best performance on most datasets. In particular, it surpasses the SwinIR [41] by even 0.48dB with $\sigma$=50 on the Urban100 dataset. We also give a visual comparison in Fig. 7. Thanks to the global receptive field, our MambaIR can achieve better structure preservation, leading to clearer edges and natural shapes.
@@
Real Image Denoising. We further turn to the real image denoising task to evaluate the robustness of MambaIR when facing real-world degradation. Following [72], we adopt the progressive training strategy for fair comparison. The results, shown in Tab. 6, suggest that our method achieves comparable performance with existing state-of-the-art models Restormer [69] and outperforms other methods such as Uformer [69] by 0.12dB PSNR on SIDD dataset, indicating the ability of our method in real image denoising.
Table 6: Quantitative comparison on the real image denosing task.